OpenStudy (anonymous):

Solve 27-6(x-10)>-3x+26 Write your solution using set builder notation.

OpenStudy (anonymous):

\[x<\frac{61}{3} \]

OpenStudy (anonymous):

Start by subtracting 27 from both sides.. \[\implies -6(x-10) \gt -3x -1\]Distribute the -6\[\implies -6x + 60 \gt -3x - 1\]Add 6x to both sides\[\implies 60 \gt 3x -1\]Add 1 to both sides\[\implies 61 \gt 3x\]Divide by 3\[\implies x \lt \frac{61}{3}\] From here we just need to do set builder notation..

OpenStudy (anonymous):

Typically for set builder you can just put the inequality you find inside the notation: \[\{x| x \lt \frac{61}{3}\}\]

OpenStudy (anonymous):

ok i see thank you very much, now only 10 more to go! lol